Solve (2D^3+3D^2D'+4DD'^2+5D'^3)z=x^2+y
1
Expert's answer
2020-07-14T17:54:39-0400

"Solution"


"(2D3+3D2D'+4DD'2+5D'3)z=x2+y"


The PDE is a homogeneous linear differential equation,thus auxilary equation is;



"2m3+3m2+4m+5=0"

"\\therefore \u2005\u200a\\implies m_1=\\frac{1}{2} (-1 + \\frac{(-63 + 2\\sqrt{1086})^{1\/3}}{3^{2\/3}} - \\frac{5}{(3 (-63 + 2\\sqrt{1086}))^{1\/3}})\\\\"

"m_2=-\frac{1}{2}- (1 + \frac{(1+i\sqrt{3})(-63 + 2\sqrt{1086})^{1/3}}{4*3^{2/3}} + \frac{5(1-i\sqrt{3})}{4(3 (-63 + 2\sqrt{1086}))^{1/3}})\\"

"m_3=-\frac{1}{2}- (1 + \frac{(1-i\sqrt{3})(-63 + 2\sqrt{1086})^{1/3}}{4*3^{2/3}} + \frac{5(1+i\sqrt{3})}{4(3 (-63 + 2\sqrt{1086}))^{1/3}})\\"

"Hence,"


"C.F=f_1\u200b(y+m_1\u200bx)+f_2\u200b(y+m_2\u200bx)+f_3\u200b(y+m_3\u200bx)"

"Now;"

"(P.I)_1\u200b=\\frac{x^2}{2D^3+3D^2D\u2032+4DD\u2032^2+5D\u2032^3\u200b}"


"\\implies \\frac{1}{2D^3\u200b}(1+(\\frac{3D'}{2D}\u200b+2(\\frac{D\u2032\u200b}{D})^2+\\frac{5}{2}\u200b(\\frac{D'}{D})^2))^{\u22121}(x^2)=\\frac{1}{2D^3\u200b}(x^2)=\\frac{x^5\u200b}{120}"


"For (P.I)_2;"

"(P.I)_2\u200b=\\frac{y}{2D^3+3D^2D\u2032+4DD\u2032^2+5D\u2032^3\u200b}"


"\\implies \\frac{1}{5D'^3\u200b}(1+(\\frac{4D}{5D'}\u200b+\\frac{3}{5}(\\frac{D\u200b}{D'})^2+\\frac{2}{5}\u200b(\\frac{D}{D'})^2))^{\u22121}(y)=\\frac{1}{5D'^3\u200b}(y)=\\frac{y^4}{120}"


"Hence,"


"C.F+(P.I)_1\u200b+(P.I)_2\u200b"

"\\implies \u2211_{i=1}^{3}\u200bf_i\u200b(y+m_i\u200bx)+\\frac{1}{120}\u200b(x^5+y^4)"